how many grams of water are produced when 2.50 mol oxygen reacts with hydrogen?
1
Expert's answer
2012-12-24T11:23:37-0500
According to reaction below, when one mol of oxygen reacts with two mol of hydrogen, two mol of water is produced. Thus, when 2.5 mol of oxygen reacted with hydrogen, 2 * 2.5 mol = 5 mol of water is produced. The molar mass of water is 18 g/mol. One mol of water weighs 18 g. 5 mol of water weights 5 * 18 = 90 g.
O2 + 2 H2 = 2 H2O
Answer: 90 g.